# Defect liquids in a weakly imbalanced bilayer Wigner Crystal

### v1

@@ I Introduction
Recent experiments across a variety of two-dimensional electron gas (2DEG) systems have provided compelling evidence for the existence of Wigner crystal (WC) phases at low electron densities [1, 2, 3, 4, 5, 6, 7, 8, 9]. Among these the bilayer WC [2] – formed in two Coulomb-coupled equal density electronic layers separated by a distance $d$ – is an especially promising platform for investigating properties of the electron solid due to the possibility of realizing different WC lattice geometries with varying interlayer coupling strength [10, *Vilk1985, 12, 13, 14, 15, 16, 17] and the enhanced stability of the bilayer crystal to higher electron densities [18, 16, 19, *rapisarda1998].
In any realistic bilayer WC system there will be some, possibly small, density mismatch $\delta n$ between electron densities in the two layers. Denoting the layer densities $n_{1}=n+\delta n$ and $n_{2}=n$, a slight imbalance $|\delta n|\ll n$ will lead to the introduction of a small concentration of defects in the ground state of the bilayer WC (if the interlayer coupling is not too weak). Even a dilute concentration of such defects may have outsized effects on the properties of the system. For example, in the monolayer WC it was recently observed that tunneling barriers for interstitials and vacancies are small, being significantly reduced compared to tunneling barriers associated with ring-exchange processes [21, 22]. The dynamics of these defects leads to kinetic magnetism with energy scales much larger than those associated with ring-exchange [21], as well as the possiblity of a self-doping instability of the monolayer WC to a metallic crystal state [22]. If the tunneling barriers are similarly small in the bilayer WC, then the dynamics of a dilute concentration of defects in a weakly imbalanced bilayer may have important implications for the magnetism and transport properties of the system.
@@ II.1 Balanced bilayer Wigner Crystal
$H=\sum_{i}\frac{\mathbf{p}_{i}^{2}}{2m}$ $++\frac{1}{2}\sum_{i\neq j}\frac{e^{2}}{\sqrt{|\mathbf{r}_{i}-\mathbf{r}_{j}|^{2}+d_{ij}^{2}}}.$ (1)
@@
In the extreme dilute limit $r_{s}\rightarrow\infty$, the potential term dominates over the kinetic term in Eq. (1) and the phase diagram as a function of $\eta$ is completely determined by minimizing the classical electrostatic energy. It is known [10, *Vilk1985, 12, 13, 14, 15, 16, 17] that there are five different phases in this limit, as depicted in Fig. 1. In order of increasing $\eta$, the phases are: one-component hexagonal (I), staggered rectangular (II), staggered square lattice (III), staggered rhombic (IV), and staggered hexagonal (V) state. Phases I-IV are connected via sequential second-order phase transitions, while the transition from IV-V transition is first-order.
@@
where the first and second terms correspond respectively to the electrostatic and vibrational energies. The dimensionless coefficients $A_{1}$ and $A_{3/2}$ as a function of $\eta$ across the various bilayer WC geometries are shown in Fig. 2.
Figure 2: Coefficients $A_{1}$ and $A_{3/2}$ as a function of $\eta$. The vertical solid (dashed) line denotes the first(second)-order phase transition (same for Figs. 5, 6 and 7).
@@ II.2 Defect model of the imbalanced bilayer Wigner crystal
Figure 3: Schematics of the energy $\varepsilon_{\text{tot}}$ as a function of VD density $\nu_{v}$ for (a, b) VD, (c, d) ID, and (e, f) IVBS state.
With the above assumptions in mind and assuming $\delta n=n_{1}-n_{2}>0$, the ground state of an imbalanced bilayer WC is described by a state having IDs (VDs) in the first (second) layer with density $\nu_{i}$ $(\nu_{v})$, subject to the constraint $\delta n=\nu_{i}+\nu_{v}$. Due to the attraction, equal amounts of IDs and VDs form IVBSs of density $\nu_{iv}=\text{min}\{\nu_{i},\nu_{v}\}$, leaving the rest of the IDs or VDs unpaired. The total energy density $\epsilon_{\text{tot}}$ as a function of $\nu_{v}$ is shown schematically in Fig. 3. There are three possible different defect liquid ground states: $\nu_{i}=\delta n,~{}\nu_{v}=0$ (ID liquid), $\nu_{i}=0,~{}\nu_{v}=\delta n$ (VD liquid), or $\nu_{i}=\nu_{v}=\delta n/2$ (IVBS liquid).
@@
We now consider the energy density $\varepsilon_{D}$ – as opposed to energy per particle, as in Eq. (2) – of the various defect ground states to first order in $\delta n$. The energy density of the balanced bilayer per layer is denoted $\varepsilon_{0}=\epsilon(r_{s})/\pi r_{s}^{2}$ and the associated chemical potential $\mu_{0}=\partial\varepsilon_{0}/\partial n$. We have
@@
and
$\varepsilon_{iv}(n,\delta n)$ $=2\varepsilon_{0}(n+\delta n/2)+\mu_{iv}(n+\delta n/2)\delta n/2$ (6)
@@
For the IVBS liquid to be the lowest-energy state, Eqs. (8) and (9) may be combined into the single condition $E^{b}_{iv}+|\Delta\varepsilon_{s}|<0$. If $E^{b}_{iv}+|\Delta\varepsilon_{s}|>0$, then the ground state is an ID liquid for $\Delta\varepsilon_{s}>0$, while for $\Delta\varepsilon_{s}<0$ the groud state is a VD liquid.
@@ II.3 Numerical procedure
Our numerical calculations employ a supercell commensurate with the defect-free bilayer WC lattice for a given $\eta$. That is, if the lattice primitive vectors are $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$, then we choose a supercell with primitive vectors $\mathbf{c}_{1}=L\mathbf{a}_{1}$ and $\mathbf{c}_{2}=L\mathbf{a}_{2}$ with $L$ an integer. A defect-free bilayer WC cell thus contains $N=2L^{2}$ electrons.
@@ III.2 Phase diagram
Figure 7: The phase diagram of the defect liquid, where red, green and blue regions denote the IVBS, VD and ID respectively. We interpolate the data in Fig. 5 to obtain a smooth phase boundary.
The defect phase diagram at finite $r_{s}$ is obtained by taking into account both the classical electrostatic and zero-point vibrational energies. Our results are shown in Fig. 7. We only show the solid phase; the solid-liquid transition occuring for $20\lesssim r_{s}(\eta)\lesssim 50$ [19, 20] is not shown. Quantum fluctuations either enhance or suppress the binding energy $E^{b}_{iv}$ and the chemical potentials $\mu_{i}$, $\mu_{v}$, thereby leading to diverse phases at finite $r_{s}$. We find the possiblity of a VD to IVBS transition with increasing density (decreasing $r_{s}$) near the II-III phase boundary, as well as an IVBS to VD transition with increasing density for $\eta\lesssim 0.2$. The kink in the phase boundary near the II to III structural transition can be traced back to the anomalous behavior of $B^{D}_{3/2}$ (see Fig. 5).
@@ IV Conclusion
In this paper we have numerically investigated the energetics of individual and paired point defects in a weakly imbalanced bilayer WC. Utilizing a simplified defect model, we have mapped out the ground state defect phase diagram as a function of interlayer distance and electron density in the semiclassical approximation. We have found three possible types of defect liquid states – intersitial defect (ID) liquid, vacancy defect (VD) liquid, and interstitial-vacancy bound state (IVBS) liquid – all of which exist in a wide range of parameters (see Fig. 7).
Tunneling of these defects has a number of consequences for the transport and magnetic properties of the weakly imbalanced bilayer WC. While the perfect bilayer WC is an insulator, the charged ID and VD in the imbalanced system will conduct current when a voltage is applied to both layers (so long as the system is sufficiently clean that the defects are not localized). The IVBS is charge-neutral and would not respond to a uniform field, but could be detected by measuring drag resistivity [35].
@@ Appendix B Finite-size scaling of the chemical potential coefficients $B^{D}$
In Figs. 5 and 5 of the main text we report the coefficeints $B_{1}^{(D)}$ and $B_{3/2}^{(D)}$ extrapolated to the thermodynamic limit $L\to\infty$. (We recall the supercell lattice vectors are taken to be $\mathbf{c}_{1,2}=L\mathbf{a}_{1,2}$, where $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$ are the bilayer WC primitive vectors). As explained in the main text, we have utilized the scaling forms $\Delta B_{1}^{D}\sim L^{-2}$ and $\Delta B^{(D)}_{3/2}\sim L^{-3}$ for the finite-size corrections, as proposed by Cockayne and Elser [30]. In Fig. 8 we show representative examples of this scaling.

### v2

@@ I Introduction
Recent experiments across a variety of two-dimensional electron gas (2DEG) systems have provided compelling evidence for the existence of Wigner crystal (WC) phases at low electron densities [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Among these the bilayer WC [2, 11, 12, 13, 14] – formed in two Coulomb-coupled equal density electronic layers separated by a distance $d$ – is an especially promising platform for investigating properties of the electron solid due to the possibility of realizing different WC lattice geometries with varying interlayer coupling strength [15, *Vilk1985, 17, 18, 19, 20, 21, 22, 23, *rapisarda1998] and the enhanced stability of the bilayer crystal to higher electron densities [25, 21, 23, *rapisarda1998].
In any realistic bilayer WC system there will be some, possibly small, density mismatch $\delta n$ between electron densities in the two layers. Denoting the layer densities $n_{1}=n+\delta n$ and $n_{2}=n$, a slight imbalance $|\delta n|\ll n$ will lead to the introduction of a small concentration of defects in the ground state of the bilayer WC (if the interlayer coupling is not too weak), as shown in Fig. 1. Even a dilute concentration of such defects may have outsized effects on the properties of the system. For example, in the monolayer WC it was recently found that tunneling barriers for interstitials and vacancies are small, being significantly reduced compared to tunneling barriers associated with ring-exchange processes [26, 27]. The dynamics of these defects leads to kinetic magnetism with energy scales much larger than those associated with ring-exchange [26], as well as the possiblity of a self-doping instability of the monolayer WC to a metallic crystal state [27]. If the tunneling barriers are similarly small in the bilayer WC, then the dynamics of a dilute concentration of defects in a weakly imbalanced bilayer may have important implications for the magnetism and transport properties of the system.
@@
Figure 1: Three types of defects – interstitial defect (ID), vacancy defect (VD), and interstitial-vacancy bound state (IVBS) – in a imbalanced bilayer Wigner crystal. Lighter points indicate projection of electron positions from the top layer onto the bottom layer.
@@ II.1 Balanced bilayer Wigner Crystal
$H=\sum_{i}\frac{\mathbf{p}_{i}^{2}}{2m}$ $+\frac{1}{2}\sum_{i\neq j}\frac{e^{2}}{\sqrt{|\mathbf{r}_{i}-\mathbf{r}_{j}|^{2}+d_{ij}^{2}}}.$ (1)
@@
In the extreme dilute limit $r_{s}\rightarrow\infty$, the potential term dominates over the kinetic term in Eq. (1) and the phase diagram as a function of $\eta$ is completely determined by minimizing the classical electrostatic energy. It is known [15, *Vilk1985, 17, 18, 19, 20, 21, 22] that there are five different phases in this limit, as depicted in Fig. 2. In order of increasing $\eta$, the phases are: one-component hexagonal (I), staggered rectangular (II), staggered square lattice (III), staggered rhombic (IV), and staggered hexagonal (V) state. Phases I-IV are connected via sequential second-order phase transitions, while the transition from IV-V transition is first-order.
@@
where the first and second terms correspond respectively to the electrostatic and vibrational energies. The dimensionless coefficients $A_{1}$ and $A_{3/2}$ can be computed numerically using the Ewald technique discussed in Sec. II.3 and Appendix A (see also Ref. [20]). Their dependencies on $\eta$ across the various bilayer WC geometries are shown in Fig. 3.
Figure 3: Coefficients $A_{1}$ and $A_{3/2}$ as a function of $\eta$. The vertical solid (dashed) line denotes the first(second)-order phase transition (same for Figs. 6, 7 and 8).
@@ II.2 Defect model of the imbalanced bilayer Wigner crystal
Figure 4: Schematic plot of the possible dependencies of the energy $\varepsilon_{\text{tot}}$ on the VD density $\nu_{v}$ for (a, b) VD, (c, d) ID, and (e, f) IVBS state.
With the above assumptions in mind and assuming $\delta n=n_{1}-n_{2}>0$, the ground state of an imbalanced bilayer WC is described by a state having IDs (VDs) in the first (second) layer with density $\nu_{i}$ $(\nu_{v})$, subject to the constraint $\delta n=\nu_{i}+\nu_{v}$. Due to the attraction, equal amounts of IDs and VDs form IVBSs of density $\nu_{iv}=\text{min}\{\nu_{i},\nu_{v}\}$, leaving the rest of the IDs or VDs unpaired. As demonstrated below, there are three possible different defect liquid ground states: $\nu_{i}=\delta n,~{}\nu_{v}=0$ (ID liquid), $\nu_{i}=0,~{}\nu_{v}=\delta n$ (VD liquid), or $\nu_{i}=\nu_{v}=\delta n/2$ (IVBS liquid).
@@
We now consider the energy density $\varepsilon_{D}$ – as opposed to energy per particle, as in Eq. (2) – of the various defect ground states to first order in $\delta n$. The energy density of the balanced bilayer per layer is denoted $\varepsilon_{0}(n)=\epsilon(r_{s})/\pi r_{s}^{2}$ where $r_{s}=1/\sqrt{\pi n}$, and the associated chemical potential $\mu_{0}=\partial\varepsilon_{0}/\partial n$. For an ID, it is straightforward to show [^1] To avoid confusion: in Eqs. (4)-(6), $n$ and $\delta n$ inside all the paratheses should be understood as the arguments of the corresponding quantities.
@@
For a VD, the system can be regarded as having vacancies with concentration $\delta n$, on top of a balanced bilayer with electrons of density $n+\delta n$ per layer. Therefore, the energy density for VD is given by
@@
Similarly, for the IVBS state, one obtains
$\varepsilon_{iv}(n,\delta n)$ $=2\varepsilon_{0}(n+\delta n/2)+\mu_{iv}(n+\delta n/2)\delta n/2$
@@
Depending on the signs of Eqs. (7)(8) and (9), the total energy density as a function of $\nu_{v}$ has different possible dependences, which are shown schematically in Fig. 4. Although we have only considered three specific cases with $\nu_{v}=0,\delta n/2,\delta n$, one can show that $\epsilon_{\text{tot}}$ changes linearly with $\nu_{v}$ when $20<\nu_{v}<\delta n/$ 2 or $\delta n/2<\nu_{v}<\delta n$ in the dilute defect limit, so these three cases are indeed the only important ones. For the IVBS liquid to be the lowest-energy state, Eqs. (8) and (9) may be combined into the single condition $E^{b}_{iv}+|\Delta\varepsilon_{s}|<0$. If $E^{b}_{iv}+|\Delta\varepsilon_{s}|>0$, then the ground state is an ID liquid for $\Delta\varepsilon_{s}>0$, while for $\Delta\varepsilon_{s}<0$ the groud state is a VD liquid.
@@ II.3 Numerical procedure
Our numerical calculations of the defect energies employ a supercell commensurate with the defect-free bilayer WC lattice for a given $\eta$. That is, if the lattice primitive vectors are $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$, then we choose a supercell with primitive vectors $\mathbf{c}_{1}=L\mathbf{a}_{1}$ and $\mathbf{c}_{2}=L\mathbf{a}_{2}$ with $L$ an integer. A defect-free bilayer WC cell thus contains $N=2L^{2}$ electrons.
@@ III.2 Phase diagram
Figure 8: The phase diagram of the defect liquid, where red, green and blue regions denote the IVBS, VD and ID respectively. We interpolate the data in Fig. 6 to obtain a smooth phase boundary. The yellow line indicates the solid-homogeneous liquid transition extracted from the data in Ref. [23] for a balanced bilayer.
The defect phase diagram at finite $r_{s}$ is obtained by taking into account both the classical electrostatic and zero-point vibrational energies. Our results are shown in Fig. 8, in which the solid-homogeneous liquid transition (yellow line) is sketched based on the results from Ref. [23, *rapisarda1998] [^3] The transition line was determined in [23, 24] for a balanced bilayer; a non-zero (but small) defect density will likely lead to correspondingly small modifications in the phase boundary.. We note that anharmonic effects beyond the harmonic approximation may become relevant at smaller $r_{s}$, which are out of the scope of this work. Quantum zero-point fluctuations either enhance or suppress the binding energy $E^{b}_{iv}$ and the chemical potentials $\mu_{i}$, $\mu_{v}$, thereby leading to diverse phases at finite $r_{s}$. We find the possiblity of a VD to IVBS transition with increasing density (decreasing $r_{s}$) near the II-III phase boundary, as well as an IVBS to VD transition with increasing density for $\eta\lesssim 0.2$. The kink in the phase boundary near the II to III structural transition can be traced back to the anomalous behavior of $B^{D}_{3/2}$ (see Fig. 6).
@@ IV Conclusion
In this paper we have numerically investigated the energetics of individual and paired point defects in a weakly imbalanced bilayer WC. Utilizing a simplified defect model, we have mapped out the ground state defect phase diagram as a function of interlayer distance and electron density in the semiclassical approximation. We have found three possible types of defect liquid states – interstitial defect (ID) liquid, vacancy defect (VD) liquid, and interstitial-vacancy bound state (IVBS) liquid – all of which exist in a wide range of parameters (see Fig. 8).
Defect tunneling process should have a number of consequences for the transport and magnetic properties of the weakly imbalanced bilayer WC. While the perfect bilayer WC is an insulator, the charged ID and VD in the imbalanced system will conduct current when a voltage is applied to both layers (so long as the system is sufficiently clean that the defects are not localized). The IVBS is charge-neutral and would not respond to a uniform field, but could be detected by measuring drag resistivity [42]. Noise spectroscopy [43] may also be a useful probe of the defect liquid phases we have described, as we expect the metallic character of the itinerant defects to lead to an enhanced current noise relative to the balanced bilayer.
@@ Appendix B Finite-size scaling of the chemical potential coefficients $B^{D}$
In Figs. 6 and 6 of the main text we report the coefficeints $B_{1}^{(D)}$ and $B_{3/2}^{(D)}$ extrapolated to the thermodynamic limit $L\to\infty$. (We recall the supercell lattice vectors are taken to be $\mathbf{c}_{1,2}=L\mathbf{a}_{1,2}$, where $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$ are the bilayer WC primitive vectors). As explained in the main text, we have utilized the scaling forms $\Delta B_{1}^{D}\sim L^{-2}$ and $\Delta B^{(D)}_{3/2}\sim L^{-3}$ for the finite-size corrections, as proposed by Cockayne and Elser [35]. In Fig. 9 we show representative examples of this scaling.